wiki:UC20170710

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Phonecon, 7/10/17

Attending: Nancy Chanover, Mark Klaene, Rachel Kuzio de Naray, Scott Anderson, Sarah Tuttle, Joanne Hughes Clark, Bill Ketzeback, Anne Verbiscer, Mary Beth Kaiser, John Wilson, John Bally


With Bruce Gillespie's retirement, we are in need of a new minute-taker for the Users Committee telecons. Anne Verbiscer kindly volunteered, although she could not do it this month due to her travel schedule. Nancy took the minutes for July with backup from Mark.


User feedback and comments from institutional representatives:

  • Colorado - John B. had nothing new to report.
  • Georgia State - Rachel had nothing new to report.
  • JHU - Mary Beth had nothing new to report.
  • Washington - Scott had nothing new to report.
  • NAPG - absent; no report
  • NMSU - absent; no report
  • Oklahoma - absent; no report
  • Wyoming - absent; no report
  • Seattle - Joanne mentioned that she has been fighting with some of the ARCTIC reduction routines; she is working with Karen Kinemuchi on this and will help to update the info on the wiki
  • Virginia - Anne had nothing new to report other than the good news of TripleSpec returning to service (see below for more info)

Telescope and Instruments Report:

Monsoon season has arrived at the site, and this has impacted observing as expected. SDSS shutdown starts today. Construction on the tunnel on HW 82 means that the road will be closed from 8 pm - 5 am Sunday through Thursday; the detour route adds approximately 1.5 hours of travel time, and this situation could last for 3+ months. There is a back road to APO but it involves creek crossings and its accessibility is weather-dependent. Anyone coming to the site is encouraged to contact the site to get the most up-to-date road travel information. The status of the 3.5m instruments (specifically ARCES, DIS, and Tspec) was discussed in detail. As to the telescope and instrument report, Mark’s more detailed material follows:

3.5-m Telescope and Instruments Highlights, 6/7/17 through 7/5/17

0) Overview

Some disruptions due to early monsoonal moisture (more clouds than precipitation) also some smoke from Arizona wildfires. However big story is instruments - or lack thereof. For anyone coming to the site from July through September: US82 at the tunnel will be closed Sunday through Thursday from 8 pm to 5 am . The published detour is through Mescalaro on US 70 and NM 244. This detour will add about 1 to 1 1/2 hours to your trip.

1) Telescope - Telescope was operational with no issues.

2) Instruments

The DIS blue and red cameras started to warm with high ambient temperatures; while not totally uncommon it became apparent that it wasn’t just a high ambient temperature problem, especially on the red camera. While several low/moderate risk fixes were attempted, it appears that the real problem is a blockage in the gas lines and the only fix is to warm/pump/flush the gas lines, which takes several days. We are working on some scheduling shuffling to make that happen. When the temperatures are higher the dark current is also higher, so DIS users should pay particular attention to dark current with long exposure times.

TripleSpec: The main body of the instrument has been cooled down and we are starting to perform initial checkout of instrument following repairs. This is changing a lot over the next few days so will provide an up-to-date status at the telecon (ed. note: see additional status updated below).

Agile has been operational this period.

Echelle failed from morning to afternoon 2 weeks ago. Ed, with remote assistance from one of the instrument builders (Dale Sanford, now at Lick Observatory), has been chest deep into troubleshooting ever since. They confirmed one bad PCI card and replaced it but are still seeing issues. Spare hardware that was provided and some of the documentation is out of date and not compatible. Consequently it is slow going, especially tests near the CCD. We are also working on getting Dale out to APO but that is dependent on his responsibilities at Lick, which are substantial.

NIC-FPS is operational.

ARCTIC is operational with no change

3) ARCSAT - Operation has been fairly routine.


TripleSpec status update

TripleSpec is back up and running! Russet and Mike Skrutskie observed together two nights ago and the report so far is positive: the focus looks fine, which is also supported by calibration observations, and there are no readout problems (which was the original source of the problem). They didn’t get enough time to check on the instrument's sensitivity but they are ready to do so during their next scheduled time in two nights. The slit has not moved so it is still in the default (1.1”) position, and the TripleSpec team requested that it be left there until the end of the quarter so they can gather more data on how the spectrograph is working overall. Mary Beth reported that Russet was very helpful on the engineering night during which Mary Beth acquired some TripleSpec data; she has not looked at the data yet but will do so very soon.


ARCES status update

Mark discussed this at length during his telescope/instrument report and Bill did not have much more to add. Regarding the ARCES failure, the instrument documentation was found to be pretty good, but the spare situation was poor to non-existent. Ed replaced a failed microtransmitter, purchased a PCI card for the ICC computer, and is in frequent communication with Dale Sanford, who has been very helpful but not totally accessible due to his own schedule constraints. Connie Rockosi may be coming out to APO the week of 7/24 (to work on eBOSS dewars), so we are exploring options for how best to get help from people who know the instrument. It is taking a lot of time to go through all of the diagnostics. The problem appears to be at the camera saddlebags, which means repairs have to proceed more slowly as we get closer to the CCD. Mary Beth asked if we know when ARCES will be back online? Nancy said she is waiting to hear from Dale and Connie, so we are in a holding pattern until we know what their travel plans are.


DIS status update

DIS red is cold but still on vacuum pump - it should be ready to go on the sky by next Wednesday (i.e the next time it is scheduled for use).


Q3 3.5-m scheduling:

Payback time for people who have been displaced due to the ARCES failure is being tracked by Russet, Ben, and Nancy. People can let us know if they are worried about making sure they are rescheduled appropriately, but every effort is being made to track this carefully.


AS4 Briefing and Discussion with Juna Kollmeier

Juna Kollmeier, the AS4 (After SDSS-IV) Director, joined the call to give a briefing on the current status of AS4 and discuss potential areas of collaboration. This is likely the start of a longer conversation about ways in which ARC (and the 3.5m) can collaborate in and contribute to AS4 science. The current status is that the AS4 proposal was submitted to the Sloan Foundation at end of June; the outcome will not be known until October. The AS4 team is currently in the phase of infrastructure and collaboration building. Juna briefly discussed the overarching science goals of AS4, which have broad appeal to ARC institution members. It will include operations in both the northern and southern hemispheres (at APO and LCO, respectively). The Local Volume Mapper, if it approved to go forward, would include the main LVM portion and 3 DESI-like spectrographs on multiple telescopes at APO. Juna welcomes ideas from the ARC community about participation in SDSS5 and synergies with other programs going on within ARC. They are currently working on a detailed document about the survey planning, which she expects will be completed this week, and ARC members are encouraged to look at that for more information. The most time-critical thing right now is that AS4 is offering incentives for institutions who provide seed funds ($1 = $1.6) until August 31. The biggest incentive is that early joiners will be able to play a significant role in shaping the survey and survey planning. More information is available on the futures web page, including the preliminary joining document (round 1 of incentives ended on June 1, we are now in round 2, and there will be a round 3).


TDAWG Update

For the newly formed Time Domain Astronomy Working Group, we have participation from UW (lots), CU, and OU, but we would like broader representation from ARC member institutions. If Users Committee representatives (outside of those 3 institutions) know of users at your institutions who are interested in time domain astronomy, please encourage them to volunteer by sending an email to Ben Williams! The work commitment will not be too onerous; telecons will begin in August.


Next Meeting:

We will still meet in August (on August 14), despite the fact that this may coincide with the 3.5m return to science after the summer shutdown. Agenda is still TBD but we may have a meeting with a slightly different flavor - perhaps a discussion devoted to just instrumentation, or potential involvement in AS4.


Acknowledgement of Russet's Scheduling Efforts

Nancy thanked Russet for her hard work on the scheduling front, which has required a lot of shuffling and time reassignments due to ARCES being out of commission. Russet has done an amazing job of keeping track of the plethora of schedule changes and helping users to still get useful science data in the face of these instrumentation challenges. Everyone on the call agreed - huge thanks to Russet for her scheduling efforts!


ACTION ITEMS:

open from previous meetings:

  • Users Committee members to query user interest in time-domain science using the 3.5-m telescope, and get back to Nancy and Ben as soon as possible. Status: CLOSED
  • Nancy will work with Ben and Russet on a web page that lists the telescope scheduling policies to avoid any future misunderstandings. Status: OPEN. Nancy sent out an email to the institutional schedulers, which then generated some questions that need to be cleared up. A new email and a web page outlining these policies still need to be developed.
  • Nancy will work with Bill to make sure that the Instrument News page reflects the current status information that users are looking for. Status: CLOSED. Nancy and Bill discussed this off-line, and Mark reminded the Users Committee reps that this page is an engineering page, and is not intended to be science-oriented (that is what the wiki is for).

new from this meeting:

  • Users Committee members should poll their users about interests in AS4 once we distribute the overview document.

Next meeting:

The next Users Committee phonecon will be on Monday 14 August at 1030 MDT — the agenda and other materials will be sent to committee members during the preceding week.

Note: See TracWiki for help on using the wiki.